# Root Defence Strategies: Ensuring Safety of LLM at the Decoding Level

### v2

@@ Abstract
Large language models (LLMs) have demonstrated immense utility across various industries. However, as LLMs advance, the risk of harmful outputs increases due to incorrect or malicious prompts. While current methods effectively address jailbreak risks, they share common limitations: 1) Judging harmful outputs from the prefill-level lacks utilization of the model’s decoding outputs, leading to relatively lower effectiveness and robustness. 2) Rejecting potentially harmful outputs based on a single evaluation can significantly impair the model’s helpfulness. To address the above issues, we examine LLMs’ capability to recognize harmful outputs, revealing and quantifying their proficiency in assessing the danger of previous tokens. Motivated by pilot experiment results, we design a robust defense mechanism at the decoding level. Our novel decoder-oriented, step-by-step defense architecture corrects the outputs of harmful queries directly rather than rejecting them outright. We introduce speculative decoding to enhance usability and facilitate deployment to boost safe decoding speed. Extensive experiments demonstrate that our approach improves model security without compromising reasoning speed. Notably, our method leverages the model’s ability to discern hazardous information, maintaining its helpfulness compared to existing methods.
@@ 1 Introduction
Large language models (LLMs) have advanced significantly in recent years, prompting growing attention from academia and industry to their safety implications (Weidinger et al., 2021; Achiam et al., 2023; Wu et al., 2023b). One of the primary safety concerns is jailbreaking, where malicious actors or errant inputs prompt LLMs to produce harmful or inappropriate content, effectively bypassing ethical guidelines. Many attempts have been made to address these risks. For instance, Meta has implemented several strategies in both pre-training and fine-tuning phases to improve the safety of their Llama-series models (Touvron et al., 2023; Dubey et al., 2024). Despite these efforts, some studies have reported that focusing too narrowly on safety may diminish the models’ general capability (Bai et al., 2022; Huang et al., 2024). Therefore, enhancing LLMs’ safety without compromising their utility has become a critical area of research.
Recent defense strategies against jailbreaks can be roughly categorized into two groups (as shown in Figure 1). The first group is prefill-level defense (Wu et al., 2023a; Phute et al., 2023; Zheng et al., 2024). It enhances the models’ protective capabilities by integrating additional security measures into the initial prompts (prefills) or refining their representation. However, this approach primarily depends on user inputs to detect harmful outputs, making it susceptible to rapidly advancing jailbreaking techniques. Moreover, this reliance can lead to inaccuracies in interpreting user intentions, thereby reducing the overall utility of the LLMs. Another group of methods is output-level defenses (Phute et al., 2023; Xu et al., 2024). It involves using safety filters that assess the potential harmfulness of model-generated outputs. This method focuses on the output of LLMs, potentially offering improved performance by directly addressing the content generated. However, this strategy typically involves a single evaluation point, which may result in false positives that could diminish the model’s utility by restricting benign outputs.
Figure 1: Examples of recenqt imperfect defenses and RDS. a) Prefill-level defenses fail to refuse the harmful query with $N$ harmful tokens. b) Output-level defenses judge the whole output in a single-point evaluation without consideration of the prefill. c) RDS conducts step-by-step assessments for each sampled token to enhance the security of LLMs at the decoder level.
@@ 2.1 LLMs’ Discriminative Capability of Decoding
To address these limitations, we explore whether LLMs can discriminate harmful from benign content during decoding, which encompasses both the prefill and the model’s generated outputs. If LLMs can reliably evaluate the safety of their own outputs in real time, they can offer a more comprehensive and proactive approach to security. Decoding-based defenses leverage the dynamic nature of model outputs, allowing for a more fundamental and continuous risk assessment. We use the hidden states of the harmful and benign queries at the top layer of the model for classifier training. Details of the classifier’s training objective is provided as follows.
@@
where $\mathbf{u}\in\mathbb{R}^{d}$ is the mean value of all hidden states of queries, $\mathbf{V}\in\mathbb{R}^{d\times m}$ represents the $m$ principal components, $\mathbf{W}\in\mathbb{R}^{1\times d}$ and $\mathbf{b}\in\mathbb{R}^{1}$ are the trainable parameters. $\hat{y}_{i}$ and $y_{i}$ represent the predicted score and the label of query, respectively.
@@ 2.2 Preliminary settings
We utilize Principal Component Analysis (PCA) to visualize the hidden states during the decoding process. To facilitate classifier training, we curate the training dataset Custom from DRO Zheng et al. (2024) to fit the classifier, consisting of 100 harmful and 100 benign queries. The evaluated LLMs are: Llama-2-chat-7B (Touvron et al., 2023), Llama-3-8b-Instruct (AI@Meta, 2024), Qwen2-7B-Instruct (Yang et al., 2024), Vicuna-7B-v1.3, and Vicuna-13B-v1.3 (Chiang et al., 2023). Notably, some models, such as Llama-2-chat-7B, have been aligned in safety.
We visualize the hidden state from the top layer of each generated token to verify the classifier ability at decoding. The outputs of harmful queries are assessed using Llama-guard (Bhatt et al., 2023), which is a safety classification model based on LLaMA-2 (Touvron et al., 2023). If refusal strings are identified in the output, it is categorized as a refusal output; otherwise, it is not. A compliant answer is assigned an evaluation score $s$ of 1, otherwise 0. The compliant outputs to harmful queries are treated as harmful outputs. Others including the refusal outputs to harmful queries and benign queries, and compliant outputs to benign queries are treated as benign outputs. In the preliminary experiment, we sample one output for each query. The initial defense of these five LLMs is presented in Appendix C.
@@ 4.1 Experimental setup
We select five defense methods as the baselines. Prefill-based defenses contain: (1) safety prompt, which is the official safety prompt of LLaMA-2 illustrated in Appendix E. The safety prompt serves as the system prompt of LLMs. (2) Self-Reminder (Wu et al., 2023a), which encapsulates the user’s query in a system prompt to remind LLMs to respond responsibly. (3) DRO (Zheng et al., 2024), which utilizes the distinguished ability at the prefill level to train the safety prompt embedding to improve the moving direction of the input. Output-based defenses contain: (4) Self-Examination (Phute et al., 2023), which checks the output by the LLM itself and filter out harmful output. (5) SafeDecoding (Xu et al., 2024), which amplifies the sampling probabilities of the output that matches the string of safety disclaimers learned from an additional trained export model.
@@ 4.2 Main Results
Secondly, RDS conducts security defense without increasing the rejection rate. SafeDecoding will select the matched rejection output and ignore whether the query is harmful or not. Therefore, SafeDecoding tends to reject benign query. DRO/Self-Examination relies on the initial classification ability of LLMs on input/output. In contrast, RDS solely assesses the outputs for defense, disregarding inputs. Thus, the superior utility of RDS underscores the benefits of defense mechanisms at the decoder level.
@@
Table 2: Evaluation results on Just-Eval. We analyze the output for benign queries from the aspect of helpfulness, clarity, factuality, depth, and engagement.
Model Defense Helpfulness Clarity Factuality Depth Engagement Average
@@ 4.5 Case study
5 Related Work
5.1 Existing Defenses
5.2 Speculative Decoding

### v3

@@ Abstract
Large language models (LLMs) have demonstrated immense utility across various industries. However, as LLMs advance, the risk of harmful outputs increases due to incorrect or malicious prompts. While current methods effectively address jailbreak risks, they share common limitations: 1) Judging harmful outputs from the prefill-level lacks utilization of the model’s decoding outputs, leading to relatively lower effectiveness and robustness. 2) Rejecting potentially harmful outputs based on a single evaluation can significantly impair the model’s helpfulness. To address the above issues, we examine LLMs’ capability to recognize harmful outputs, revealing and quantifying their proficiency in assessing the danger of previous tokens. Motivated by pilot experiment results, we design a robust defense mechanism at the decoding level. Our novel decoder-oriented, step-by-step defense architecture corrects the outputs of harmful queries directly rather than rejecting them outright. We introduce speculative decoding to enhance usability and facilitate deployment to boost safe decoding speed. Extensive experiments demonstrate that our method improves model security without compromising reasoning speed. Notably, our method leverages the model’s ability to discern hazardous information, maintaining its helpfulness compared to existing methods [^1] Our code is publicly available at:https://github.com/zengxy20/RDS.
@@ 1 Introduction
Large language models (LLMs) have advanced significantly in recent years, prompting growing attention from academia and industry to their safety implications (Weidinger et al., 2021; Achiam et al., 2023; Wu et al., 2023). One of the primary safety concerns is jailbreaking, where malicious actors or errant inputs prompt LLMs to produce harmful or inappropriate content, effectively bypassing ethical guidelines. Many attempts have been made to address these risks. For instance, Meta has implemented several strategies in both pre-training and fine-tuning phases to improve the safety of their Llama-series models (Touvron et al., 2023; Dubey et al., 2024). Despite these efforts, some studies have reported that focusing too narrowly on safety may diminish the models’ general capability (Bai et al., 2022; Huang et al., 2024). Therefore, enhancing LLMs’ safety without compromising their utility has become a critical area of research.
Recent defense strategies against jailbreaks can be roughly categorized into two groups (as shown in Figure 1). The first group is prefill-level defense (Xie et al., 2023; Phute et al., 2023; Zheng et al., 2024). It enhances the models’ protective capabilities by integrating additional security measures into the initial prompts (prefills) or refining their representation. However, this approach primarily depends on user inputs to detect harmful outputs, making it susceptible to rapidly advancing jailbreaking techniques. Moreover, this reliance can lead to inaccuracies in interpreting user intentions, thereby reducing the overall utility of the LLMs. Another group of methods is output-level defenses (Phute et al., 2023; Xu et al., 2024). It involves using safety filters that assess the potential harmfulness of model-generated outputs. This method focuses on the output of LLMs, potentially offering improved performance by directly addressing the content generated. However, this strategy typically involves a single evaluation point, which may result in false positives that could diminish the model’s utility by restricting benign outputs.
Figure 1: Examples of recent imperfect defenses and RDS. a) Prefill-level defenses fail to refuse the harmful query with $N$ harmful tokens. b) Output-level defenses judge the whole output in a single-point evaluation without consideration of the prefill. c) RDS conducts step-by-step assessments for each sampled token to enhance the security of LLMs at the decoder level.
@@
2 Related Work
2.1 Existing Defenses
2.2 Jailbreak Attacks
Jailbreak attacks target the security mechanisms of LLMs with the objective of circumventing them to generate unauthorized content. These attacks pose risks of privacy breaches, intellectual property theft, and misuse of model services.
Previous studies (Liu et al., 2023; Wei et al., 2024) focus on prompt engineering as a means to compromise the security of LLMs effectively. Alternative approaches employ feature-level attacks to implicitly alter the internal architecture of LLMs (Guo et al., 2024; Wang et al., 2024). For instance, GCG (Zou et al., 2023) combines greedy with gradient-based search techniques to generate universal adversarial suffixes. After concatenated the suffixes to the queries, LLMs will answer the harmful queries previously refused to answer.
2.3 Speculative Decoding
@@ 3.1 LLMs’ Discriminative Capability of Decoding
To address these limitations, we explore whether LLMs can discriminate harmful from benign content during decoding, which encompasses both the prefill and the model’s generated outputs. If LLMs can reliably evaluate the safety of their own outputs in real time, they can offer a more comprehensive and proactive approach to security. Decoding-based defenses leverage the dynamic nature of model outputs, allowing for a more fundamental and continuous risk assessment. We use the hidden states of the harmful and benign queries from Custom (Zheng et al., 2024) at the top layer of the model for classifier training. Details of the classifier’s training objective is provided as follows.
@@
where $\mathbf{u}\in\mathbb{R}^{d}$ is the mean value of all hidden states of queries, $\mathbf{V}\in\mathbb{R}^{d\times m}$ represents the $m$ principal components, $\mathbf{W}\in\mathbb{R}^{1\times d}$ and $\mathbf{b}\in\mathbb{R}^{1}$ are the trainable parameters. $\hat{y}_{i}$ and $y_{i}$ represent the predicted score and the label of query, respectively. For harmful queries, $y_{i}=1$, while for benign queries, $y_{i}=0$.
@@ 3.2 Preliminary settings
We utilize Principal Component Analysis (PCA) to visualize the hidden states during the decoding process. To facilitate classifier training, we curate the training dataset Custom from DRO Zheng et al. (2024) to fit the classifier, consisting of 100 harmful and 100 benign queries. The evaluated LLMs are: Llama-2-chat-7B (Touvron et al., 2023), Llama-3-8b-Instruct (Grattafiori et al., 2024), Qwen2-7B-Instruct (Yang et al., 2024), Vicuna-7B-v1.3, and Vicuna-13B-v1.3 (Chiang et al., 2023). Notably, some models, such as Llama-2-chat-7B, have been aligned in safety.
We visualize the hidden state from the top layer of each generated token to verify the classifier ability at decoding. The outputs of harmful queries are assessed using Llama-guard (Bhatt et al., 2023), which is a safety classification model based on LLaMA-2 (Touvron et al., 2023). While the output of benign queries are evaluated through string matching with refusal modules. If refusal strings are identified in the output, it is categorized as a refusal output; otherwise, it is not. A compliant answer is assigned an evaluation score $s$ of 1, otherwise 0. The compliant outputs to harmful queries are treated as harmful outputs. Others including the refusal outputs to harmful queries and benign queries, and compliant outputs to benign queries are treated as benign outputs. In the preliminary experiment, we sample one output for each query. The initial defense of these five LLMs is presented in Appendix C.
@@ 5.1 Experimental setup
We select five defense methods as the baselines. Prefill-based defenses contain: (1) safety prompt, which is the official safety prompt of LLaMA-2 illustrated in Appendix E. The safety prompt serves as the system prompt of LLMs. (2) Self-Reminder (Xie et al., 2023), which encapsulates the user’s query in a system prompt to remind LLMs to respond responsibly. (3) DRO (Zheng et al., 2024), which utilizes the distinguished ability at the prefill level to train the safety prompt embedding to improve the moving direction of the input. Output-based defenses contain: (4) Self-Examination (Phute et al., 2023), which checks the output by the LLM itself and filter out harmful output. (5) SafeDecoding (Xu et al., 2024), which amplifies the sampling probabilities of the output that matches the string of safety disclaimers learned from an additional trained export model.
@@ 5.2 Main Results
Secondly, RDS conducts security defense without increasing the rejection rate. Although some methods perform well on defense, there is a catastrophic refusal phenomenon in terms of their impact on the effectiveness of LLMs, i.e., 24.0 (SafeDecoding) vs. 2.0 (No defense) on Vicuna-7B, 38.0 (SafeDecoding) and 36.0 (DRO) vs. 10.0 (No defense) on Vicuna-13B. RDS shows fewer refusal results compared to the existing security defenses, i.e., 6.0 (RDS) vs. 38.0 (SafeDecoding) on Vicuna-13B, 32.5 (RDS) vs. 100 (Self-Examination) on Llma2. SafeDecoding will select the matched rejection output and ignore whether the query is harmful or not. Therefore, SafeDecoding tends to reject benign query. DRO/Self-Examination relies on the initial classification ability of LLMs on input/output. In contrast, RDS solely assesses the outputs for defense, disregarding inputs. Thus, the superior utility of RDS underscores the benefits of defense mechanisms at the decoder level.
We then conduct an ablation study to investigate the impact of the two primary modules of RDS: the classifier and speculative head. ”-w/o SD” removes the speculative head in RDS and follows the original multi-layer blocks to compute the hidden state. ”No defense” can be seen as a variant that eliminates both the classifier and speculative head. Table 1 demonstrates that the classifier plays a crucial role in filtering out harmful tokens, while the speculative head has little impact on safety improvement. Table 3 highlights the contribution of the speculative head to the inference speed, which is consistent with our design.
@@
-w/o SD 0 0 0 0 0 12 6.0
Table 2: Evaluation results on Just-Eval. We analyze the output for benign queries from the aspect of helpfulness (H), clarity (C), factuality (F), depth (D), and engagement (E).
Model Defense H C F D E Average